4500+[4400/(1+R*)]=3500+[5500/(1+R*)]

4500–3500=[5500/(1+R*)]–[4400/(1+R*)]

1000 = (5500 – 4400) / (1+R*)
1000 = 1100 / (1+R*)

R*= 0.1 or R*= 10%

Can you explain the last two parts in this exercise?
Answer #1

We start with the given equation

4500 + [4400/(1+R*)] = 3500+ [5500/(1+R*)]

4500 –3500 = [5500/(1+R*)] – [4400/(1+R*)]

1000 = (5500 – 4400) / (1+R*)

1000 / 1 = 1100 / (1+R*)

1000(1 + R*) = 1100 ........... (Cross multiply)

1000 + 1000R* = 1100

1000R* = 1100 - 1000

1000R* = 100

R* = 100/1000 = 0.1

This gives R* = 0.1 or 0.1 x 100 = 10%.